OpenStudy (anonymous):

Find the future value and interest earned if $56,780 is invested at 5.3% compounded. (a) quaterly for 23 quaters (b) continously for 15 yrs.

OpenStudy (anonymous):

Can anyone help me??

OpenStudy (anonymous):

is the 5.3% aer, gross, interest per month per year ... need more information i'm afraid

OpenStudy (radar):

Use the compound interest formula. I will try and put it here. A=P(1+r/n)^nt Where A= balance P = Principal r= annual interest rate n=compounding time in years

OpenStudy (radar):

For part b. use the interest formula for continious compounding which is \[A=Pe ^{rt}\] the e is approximately 271828 it is called the "natural base" it is used in many exponential problems. Do you need further help?

OpenStudy (radar):

ooop I omitted the decimal point for the value of e. It is 2.71828

OpenStudy (radar):

Part a. The equation would appear like this: 56780(1+.053/4)^(4)(5.75) The 5.75 is the 23 quarters converted to years. 56780(1.01325)^23=$76,855.95

OpenStudy (radar):

part b, I leave for you to try.